# Asymptotic theory of in-context learning by linear attention

### v1

@@ Acknowledgements
The derivation of this result closely follows that of Result 6. We analyze the cases of $\tau<1$ and $\tau>1$ separately. For $\tau<1$, the equation in (SI.5.6) simplifies to (SI.5.41) as $\lambda\to 0^{+}$. For $\tau>1$, $\xi$ approaches zero as $\lambda\to 0^{+}$. Substituting these estimates into (SI.5.45) then yields (SI.5.52) after some detailed calculations.
Appendix A Equivalent Statistical Representations
@@
In this appendix, we present an equivalent (but simplified) statistical model for the regression vector $H_{Z}$ defined in (16). This statistically-equivalent model will simplify the moment calculations in Section SI.2 and the random matrix analysis in Section SI.4.
@@
$H_{Z}\overset{(d)}{=}(d/\ell)M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\begin{bmatrix}h^{\top}M_{w},&(a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix},$ (A.1)
@@
$(M_{w})e_{1}=\frac{w}{\lVert w\rVert},$ (A.3)
@@
$h\coloneqq\begin{bmatrix}\frac{{\theta_{\epsilon}}a}{\sqrt{d}}+\frac{a^{2}}{d}+\theta_{q}^{2}\\ \big{[}(\theta_{\epsilon}+a/\sqrt{d})^{2}+\theta_{q}^{2}\big{]}^{1/2}{g}/{\sqrt{d}}\end{bmatrix};$ (A.4)
@@
$\theta_{\epsilon}={\lVert v_{\epsilon}\rVert}/{\sqrt{d}}\qquad\text{and}\qquad\theta_{q}=\lVert q\rVert/\sqrt{d}.$ (A.5)
@@
For two random variables $A$ and $B$, the notation $A\overset{(d)}{=}B$ indicates that $A$ and $B$ have identical probability distributions. Note that $A$ and $B$ can be either scalars [as in the case of (A.2)], or matrices of matching dimensions [as in the case of (A.1)].
@@
A concrete construction of the symmetric and orthonormal matrix $M_{w}$ satisfying (A.3) can be based on the Householder transformation [54, 55, 56].
@@
$x_{\ell+1}\overset{(d)}{=}\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}s\\ u\end{bmatrix},$ (A.6)
where $s\sim\mathcal{N}(0,1)$ and $u\sim\mathcal{N}(0,I_{d-1})$ are two independent normal random variables (vectors), and $M_{w}$ is the symmetric orthonormal matrix specified in (A.3). Note that $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise. The representation in (A.2) then follows immediately from (A.6) and the identity in (A.3).
To show (A.1), we first reparameterize the $d\times\ell$ Gaussian data matrix $X$ as
$X=M_{w}\begin{bmatrix}a&q^{\top}\\ p&U\end{bmatrix}M_{v_{\epsilon}}/\sqrt{d}.$ (A.7)
@@
$M_{v_{\epsilon}}e_{1}=\frac{v_{\epsilon}}{\lVert v_{\epsilon}\rVert},$ (A.8)
where $e_{1}$ denotes the first natural basis vector in $\mathbb{R}^{\ell}$. Since the data matrix $X$, the task vector $w$, and the noise vector $v_{\epsilon}$ are mutually independent, it is straightforward to verify via the rotational symmetry of the isotropic normal distribution that both sides of (A.7) have identical probability distributions. Using this new representation, we have
$Xv_{\epsilon}=\theta_{\epsilon}M_{w}\begin{bmatrix}a\\ p\end{bmatrix}.$ (A.9)
@@
$X^{\top}w=M_{v_{\epsilon}}\begin{bmatrix}a\\ q\end{bmatrix},$ (A.10)
@@
$XX^{\top}w=\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}a^{2}+\lVert q\rVert^{2}\\ ap+Uq\end{bmatrix}.$ (A.11)
Combining (A.10) and (A.11) yields
$Xy$ $=XX^{\top}w+Xv_{\epsilon}$ (A.12)
$=M_{w}\begin{bmatrix}\theta_{\epsilon}a+a^{2}/\sqrt{d}+\theta_{q}^{2}\sqrt{d}\\ (\theta_{\epsilon}+a/\sqrt{d})p+Uq/\sqrt{d}\end{bmatrix}.$ (A.13)
@@
$\frac{1}{\sqrt{d}}Xy\overset{(d)}{=}M_{w}h,$ (A.14)
where $h$ is the random vector defined in (A.4).
Lastly, we consider the term $y^{\top}y$ in (16). Since $y=X^{\top}w+v_{\epsilon}$,
$y^{\top}y$ $=\lVert X^{\top}w+v_{\epsilon}\rVert^{2}$ (A.15)
$=\lVert X^{\top}w+\theta_{\epsilon}\sqrt{d}M_{v_{\epsilon}}e_{1}\rVert^{2}$ (A.16)
$=(a+\theta_{\epsilon}\sqrt{d})^{2}+\theta_{q}^{2}d,$ (A.17)
where the second equality follows from (A.8) and to reach the last equality we have used the representation in (A.10). To show (A.1), we recall the definition of $H_{Z}$ in (16). Substituting (A.6), (A.14) and (A.17) into (16), we are done. ∎
Appendix B The Stieltjes Transforms of Wishart Ensembles
@@
$\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\mathcal{M}_{\kappa}(\nu)$ (B.2)
@@
$\mathcal{M}_{\kappa}(\nu)\coloneqq\frac{2}{\nu+1-1/\kappa+\left[(\nu+1-1/\kappa)^{2}+4\nu/\kappa\right]^{1/2}}.$ (B.3)
@@
$\frac{1}{\mathcal{M}_{\kappa}(\nu)}=\frac{1}{1+\mathcal{M}_{\kappa}(\nu)/\kappa}+\nu.$ (B.4)
@@
$\frac{1}{d}\operatorname{tr}F^{2}(\nu)\simeq-\mathcal{M}^{\prime}_{\kappa}(\nu)=\frac{\mathcal{M}^{2}_{\kappa}(\nu)}{1-\frac{\kappa\mathcal{M}_{\kappa}^{2}(\nu)}{[\kappa+\mathcal{M}_{\kappa}(\nu)]^{2}}}.$ (B.5)
For the remainder of this appendix, we will further explore the self-consistent equation given by (SI.4.7). We will show that the solution $\chi_{\pi}$ and its derivative $\frac{d}{d\pi}\chi_{\pi}$, at $\pi=0$, can be characterized by the function $\mathcal{M}_{\kappa}(\nu)$ in (B.3). To start, note that at $\pi=0$, the equation in (SI.4.7) can be written as
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}.$ (B.6)
@@
$(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I_{d+1})^{-1}=\begin{bmatrix}F_{R}(\nu_{0})+a^{\ast}(1+\rho)^{2}F_{R}(\nu_{0})b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&-a^{\ast}(1+\rho)F_{R}(\nu_{0})b_{\mathrm{tr}}\\ -a^{\ast}(1+\rho)b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&a^{\ast}\end{bmatrix},$ (B.7)
where $F_{R}(\cdot)$ is the function defined in (B.1),
$\nu_{0}=\frac{1+\rho}{\alpha}+\lambda(1+\chi_{0})$ (B.8)
@@
$\frac{1}{a^{\ast}}=(1+\rho)^{2}+\lambda(1+\chi_{0})-(1+\rho)^{2}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}.$ (B.9)
From (B.7), the equation (B.6) becomes
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}F_{R}(\nu_{0})-(1+\rho)^{2}\frac{a^{\ast}\lambda(1+\chi_{0})}{d}\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}.$ (B.10)
@@
$b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}\leq 1\qquad\text{and}\qquad\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}\leq\frac{1}{\nu_{0}}\leq\frac{\alpha}{1+\rho}.$ (B.11)
Substituting the first inequality above into (B.9) gives us
$a^{\ast}\lambda(1+\chi_{0})\leq 1.$ (B.12)
Combining this estimate with the second inequality in (B.11), we can conclude that the last term on the right-hand side of (B.10) is negligible as $d\to\infty$. Moreover, using the asymptotic characterization given in (B.2), the equation (B.10) leads to
$\frac{\tau\chi_{0}}{1+\chi_{0}}\simeq 1-\lambda(1+\chi_{0})\mathcal{M}_{\kappa}(\nu_{0}).$ (B.13)
@@
$\xi_{0}=\lambda(1+\chi_{0}),$ (B.14)
and also recalling the definition of $\nu_{0}$ in (B.8), we can further transform (B.13) to
$\xi_{0}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi_{0}\right)-\frac{\tau\lambda}{\xi_{0}}\simeq 1-\tau.$ (B.15)

### v2

@@ Acknowledgments
The derivation of this result closely follows that of Result 6. We analyze the cases of $\tau<1$ and $\tau>1$ separately. For $\tau<1$, the equation in (SI.5.6) simplifies to (SI.5.41) as $\lambda\to 0^{+}$. For $\tau>1$, $\xi$ approaches zero as $\lambda\to 0^{+}$. Substituting these estimates into (SI.5.45) then yields (SI.5.51) after some detailed calculations.
SI.5.4 A note on ordering of $\lambda\to 0^{+}$ and $\alpha\to\infty$ limits
To understand why the divergence in $\alpha$ occurs for certain configurations of $\kappa$ and $\tau$, we investigate the problem where $\alpha$ is taken to infinity before $\lambda\to 0^{+}$, which is equivalent to first taking $\ell$ to infinity. This means that our features $H_{Z}$ have no token $x_{i}$ or noise $\eta$ disorder remaining, and instead depend only on $w$ and query $x_{\ell+1}$ as
$H_{Z}=\begin{bmatrix}x_{\ell+1}w^{\top}&\frac{1}{d}\left(\|w\|^{2}+\rho\right)x_{\ell+1}\end{bmatrix}$ (SI.5.52)
Computing the ridgeless limit of the generalisation error for the finite $n$ predictor (21) on the above $H_{Z}$ gives the following ridgeless, infinite-$\alpha$ generalisation error
$\lim_{\lambda\to 0^{+}}$ $\lim_{\alpha\to\infty}e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)=\begin{dcases}1-\kappa+\rho+\frac{\rho\kappa^{2}}{(\tau-\kappa)(1-\kappa)}&\kappa\leq\text{min}(\tau,1)\\ 1-\tau+\rho+\rho\frac{\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau\\ \rho+\frac{\rho\kappa}{(\tau-1)(\kappa-1)}&\tau>1,\kappa>1\end{dcases},$ (SI.5.53)
noting that the second and third cases match their equivalent finite cases in (35) in the manuscript. This result follows from applying the asymptotics given in prior works for the generalization error of ridgeless regression with structured covariates [38, 39, 41]. By comparison with (35), we see that switching the order of limits in $\alpha$ and $\lambda$ has mitigated the divergence for $\kappa\leq\min(\tau,1)$.
As these limits do not commute, which order is correct? Simulations with arbitrarily small numerical ridge parameter at large $\alpha$ track the (35) curve, not the (SI.5.53) curve. These numerical experiments are computed with the ridge scaling $n/d$ that we use elsewhere. Were we to scale this regularisation with some power of $\alpha$, we could remove this large $\alpha$ divergence; whether or not that choice is either theoretically or practically motivated is a question we leave open for future work.
SI.6 Task Generalization and Bayesian Estimators
In the main document, we have compared the performance of our linear transformer with that of two Bayesian estimators: the discrete task prior estimator and the optimal ridge estimator. Here we will derive various claims about these estimators made in the main discussion, particularly relating to their behavior against task diversity $\kappa$.
SI.6.1 Ridge estimator error
Here we prove claim that the ridge estimator error is characterized by
$e^{\text{ridge}}_{\text{ICL}}=\rho\left(1+\frac{1}{\alpha}\mathcal{M}_{\alpha}\left(\frac{\rho}{\alpha}\right)\right)=e^{\text{ridge}}_{\text{IDG}}\,.$
First, it is obvious that we must have $e^{\text{ridge}}_{\text{ICL}}=e^{\text{ridge}}_{\text{IDG}}$, as a ridge estimator by definition only operates on a single context (i.e. a single task); it cannot distinguish between $w$ drawn from $\mathcal{N}(0,I_{d})$ or $w$ drawn uniformly from $\{w_{1},\cdots,w_{k}\}$ each drawn from $\mathcal{N}(0,I_{d})$. Thus in the following we will not distinguish between these two testing regimes, and refer only to task $w\sim\mathcal{N}(0,I_{d})$.
Let $x_{1},\ldots,x_{\ell}$ be a collection of data vectors. We observe
$y_{i}=x_{i}^{\top}w+\epsilon_{i},\quad\text{for }i=1,2,\ldots,\ell,$ (SI.6.1)
where $\epsilon_{i}\sim\mathcal{N}(0,\rho)$ is the observation noise. Let $y=[y_{i}]_{i\leq\ell}\in\mathbb{R}^{\ell}$. Then we can write down second moments as
$\mathbb{E}[ww^{\top}]=I_{d},\qquad\mathbb{E}[yw^{\top}]=X^{\top},\qquad\text{and }\mathbb{E}[yy^{\top}]=X^{\top}X+\rho I_{\ell},$ (SI.6.2)
where $X=[x_{1},x_{2},\ldots,x_{\ell}]$ is the data matrix in $\mathbb{R}^{d\times\ell}$. Since $y$ and $w$ are joint Gaussian random vectors, we can write
$\beta=(XX^{\top}+\rho I_{d})^{-1}Xy+Z,$ (SI.6.3)
where $Z$ is a centered Gaussian vector independent of $y$. Note that the matrix in front of $y$ in the above expression is chosen so that the cross-correlation $\mathbb{E}[yw^{\top}]$ matches that in (SI.6.2). It then immediately follows that the Bayesian estimator of $w$ is
$w^{\text{ridge}}_{\text{Bayes}}=(XX^{\top}+\rho I_{d})^{-1}Xy.$ (SI.6.4)
Given a new input vector $x_{\ell+1}$ and its label $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon_{\ell+1}$, the Bayesian optimal estimator of the new label is
$\hat{y}_{\ell+1}$ $=x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}},$ (SI.6.5)
and the corresponding Bayesian prediction error is
$e_{\text{Bayesian}}$ $=\mathbb{E}[(y_{\ell+1}-x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}})]$ (SI.6.6)
$=\rho+\frac{1}{d}\mathbb{E}\left[\lVert w-w^{\text{ridge}}_{\text{Bayes}}\rVert^{2}\right]$ (SI.6.7)
$\simeq\rho\left(1+\frac{1}{\alpha}\mathcal{M}_{\alpha}\left(\frac{\rho}{\alpha(1-\gamma)}\right)\right),$ (SI.6.8)
where $\mathcal{M}_{\alpha}(x)$ is the function defined in (SI.8.4). In the large $\alpha$ limit, we have
$e_{\text{Bayesian}}\simeq\rho(1+1/\alpha).$ (SI.6.9)
SI.6.2 Asymptotics of the linear transformer in $\kappa$
We wish to expand (SI.5.40) and (SI.5.51) in large $\kappa$ to study the decay of $g_{\text{task}}=e_{\text{ICL}}-e_{\text{IDG}}$ toward 0 as $\kappa\to\infty$. By expanding auxillary variables in $\kappa$, we have
$\mathcal{M}_{\kappa}(q)\approx\frac{1}{1+q}\left(1+\frac{1}{(1+q)^{2}\kappa}\right)\,,\qquad\mathcal{M}^{\prime}_{\kappa}(q)\approx-\frac{1}{(1+q)^{2}}\left(1+\frac{3}{(1+q)^{2}\kappa}\right)$
and
$\mu=q\mathcal{M}_{\kappa/\tau}(q)\approx\frac{q}{1+q}\left(1+\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad\xi\approx\frac{1-\tau}{\tau}(1+q)\left(1-\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad p\approx 1+\frac{\tau^{2}}{(1+\tau)^{2}\kappa}\,.$
Substituting these simplifications into (SI.5.40) and (SI.5.51) gives the desired characterization
$g^{\text{LT}}_{\text{task}}=0+\frac{1}{\kappa}\begin{cases}c_{1}&\tau<1\\ c_{2}&\tau>1\\ \end{cases}+\mathcal{O}\left(\frac{1}{\kappa^{2}}\right)$
for
$c_{1}$ $\equiv\frac{2\left(\left(-2q-\frac{1}{2}\right)\tau^{2}+\left(q^{2}+\left(-\rho+3\right)q+1-\rho\right)\tau-\left(1+q\right)\left(q+\frac{1}{2}-\frac{\rho}{2}\right)\right)\tau}{\left(-1+\tau\right)\left(1+q\right)^{3}}\,,$ (SI.6.10)
$c_{2}$ $\equiv 2\frac{q^{2}}{(1+q)^{3}}+\frac{1}{(\tau-1)(q+1)^{2}}\left(\rho+q-\frac{q^{2}}{q+1}\right)\,.$ (SI.6.11)
SI.6.3 Asymptotics of the dMMSE estimator in $\kappa$
To study the slowness of the dMMSE estimator more explicitly, consider the $\alpha\to\infty$ limit. The exponential weight terms in the estimator in this limit behave as
$e^{-\frac{1}{2\rho}\sum_{i=1}^{\ell}\left(y_{i}-w_{j}^{\top}x_{i}\right)^{2}}\to e^{-\frac{\ell}{2\rho}\left(\frac{1}{d}\|w^{\ast}-w_{j}\|^{2}+\rho\right)}\,,$
and these weightings exponentially favor choosing $\omega_{j}$ that minimizes $\|w^{\ast}-w_{j}\|^{2}$ over the set of $k$ training tasks $w_{j}$. It’s immediately clear that $e_{\text{IDG}}^{\text{dMMSE}}=\rho$ in this limit as the minimal value of $\|w^{\ast}-w_{j}\|^{2}$ when $w^{\ast}\in\{w_{1},\cdots,w_{k}\}$ is 0. Taking
$w_{\text{est}}(w^{\ast},w_{i})=\text{argmin}_{i\in[k]}\|w^{\ast}-w_{i}\|^{2}$
we have
$g^{\text{dMMSE}}_{\text{task}}$ $=\frac{1}{d}\mathbb{E}_{w^{\ast}\sim\mathcal{P}_{\text{test}}}\left[\mathbb{E}_{w_{i}\sim\mathcal{P}_{\text{train}}}\left[\min_{i\in[k]}\|w^{\ast}-w_{i}\|^{2}\right]\right].$ (SI.6.12)
It is equivalent in high-dimensions to work with $w^{\ast},w_{i}$ sampled uniformly on the sphere $\mathcal{S}^{d-1}(\sqrt{d}).$ For analytical tractability, we will proceed with this. We exploit spherical symmetry in both $w^{\ast}$ and $w_{i}$ to simplify
$\|w^{\ast}-w_{i}\|^{2}$ $=2d-2\sqrt{d}x_{i}\qquad\text{ for $x_{i}$ a single component of vector $w_{i}\sim\text{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$}$
$=2d-2\sqrt{d}(2\sqrt{d}\,b_{i}-\sqrt{d})\qquad\text{for $b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)$}$
$=4d-4db_{i}$
This gives
$g_{\text{task}}$ $=\rho+4-4\,\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right].$
We wish to study the behavior of this expected maximum of a set of $k$ Beta variables as $k$ grows large. We are thus interested in calculating
$\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right],$ (SI.6.13)
Let $D=\frac{d-1}{2}$. Following the standard procedure for deriving the PDF for the distribution of the extreme value, have
$p(b_{i}=x)=\frac{1}{B(D)}(1-x)^{D-1}x^{D-1}$ (SI.6.14)
where
$B(D)=\frac{\Gamma(D)^{2}}{\Gamma(2D)}.$ (SI.6.15)
Then
[…]
and
[…]
and
[…]
Note that the integral can be written as
[…]
To proceed, let’s consider the variable substitution
[…]
By this definition, $t(M=0)=\infty$, $t(M=1)=0$ Plugging in
[…]
Naively, it looks like one could get a simple series expansion in $1/k$ through integration by parts, but this will fail because $\left.\frac{dM}{dt}\right|_{t=0}=\infty$ for $D>1$. One can see this by differentiating (SI.6.20) with respect to $t$. Instead, have
[…]
where we defined
$\epsilon(t/k)\equiv 1-M(t/k).$ (SI.6.23)
Now, we can expand $\epsilon(t/k)$ and integrate. We have to find an appropriate series expansion from the implicit definition of $\epsilon(t/k)$.
[…]
To get the leading behavior, one proceeds by expanding in leading order to $1-M$ or the right and to $t/k$ on the left
$1-\frac{t}{k}=1-\frac{\epsilon(t/k)^{D}}{B(D)}\quad\implies\epsilon\approx\left(\frac{B(D)t}{k}\right)^{1/D}$ (SI.6.25)
which in turn implies the leading behavior of the expectation we are looking for is
$1-\frac{B(D)^{1/D}\Gamma(1/D+1)}{k^{1/D}}$ (SI.6.26)
From this calculation we recover that the decrease of $g_{\text{task}}^{\text{dMMSE}}$ towards 0 is explicitly dimensionally cursed at $\alpha\to\infty$. We expect a similar rate to hold for finite $\alpha$ by the curse-of-dimensionality arguments given previously.
SI.7 Equivalent Statistical Representations
@@
In this appendix, we present an equivalent (but simplified) statistical model for the regression vector $H_{Z}$. This statistica lly-equivalent model will simplify the moment calculations in Section SI.2 and the random matrix analysis in Section SI.4.
@@
$H_{Z}\overset{(d)}{=}(d/\ell)M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\begin{bmatrix}h^{\top}M_{w},&(a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix},$ (SI.7.1)
@@
$(M_{w})e_{1}=\frac{w}{\lVert w\rVert},$ (SI.7.3)
@@
$h\coloneqq\begin{bmatrix}\frac{{\theta_{\epsilon}}a}{\sqrt{d}}+\frac{a^{2}}{d}+\theta_{q}^{2}\\ \big{[}(\theta_{\epsilon}+a/\sqrt{d})^{2}+\theta_{q}^{2}\big{]}^{1/2}{g}/{\sqrt{d}}\end{bmatrix};$ (SI.7.4)
@@
$\theta_{\epsilon}={\lVert v_{\epsilon}\rVert}/{\sqrt{d}}\qquad\text{and}\qquad\theta_{q}=\lVert q\rVert/\sqrt{d}.$ (SI.7.5)
@@
For two random variables $A$ and $B$, the notation $A\overset{(d)}{=}B$ indicates that $A$ and $B$ have identical probability distributions. Note that $A$ and $B$ can be either scalars [as in the case of (SI.7.2)], or matrices of matching dimensions [as in the case of (SI.7.1)].
@@
A concrete construction of the symmetric and orthonormal matrix $M_{w}$ satisfying (SI.7.3) can be based on the Householder transformation [60, 61, 62].
@@
$x_{\ell+1}\overset{(d)}{=}\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}s\\ u\end{bmatrix},$ (SI.7.6)
where $s\sim\mathcal{N}(0,1)$ and $u\sim\mathcal{N}(0,I_{d-1})$ are two independent normal random variables (vectors), and $M_{w}$ is the symmetric orthonormal matrix specified in (SI.7.3). Note that $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise. The representation in (SI.7.2) then follows immediately from (SI.7.6) and the identity in (SI.7.3).
To show (SI.7.1), we first reparameterize the $d\times\ell$ Gaussian data matrix $X$ as
$X=M_{w}\begin{bmatrix}a&q^{\top}\\ p&U\end{bmatrix}M_{v_{\epsilon}}/\sqrt{d}.$ (SI.7.7)
@@
$M_{v_{\epsilon}}e_{1}=\frac{v_{\epsilon}}{\lVert v_{\epsilon}\rVert},$ (SI.7.8)
where $e_{1}$ denotes the first natural basis vector in $\mathbb{R}^{\ell}$. Since the data matrix $X$, the task vector $w$, and the noise vector $v_{\epsilon}$ are mutually independent, it is straightforward to verify via the rotational symmetry of the isotropic normal distribution that both sides of (SI.7.7) have identical probability distributions. Using this new representation, we have
$Xv_{\epsilon}=\theta_{\epsilon}M_{w}\begin{bmatrix}a\\ p\end{bmatrix}.$ (SI.7.9)
@@
$X^{\top}w=M_{v_{\epsilon}}\begin{bmatrix}a\\ q\end{bmatrix},$ (SI.7.10)
@@
$XX^{\top}w=\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}a^{2}+\lVert q\rVert^{2}\\ ap+Uq\end{bmatrix}.$ (SI.7.11)
Combining (SI.7.10) and (SI.7.11) yields
$Xy$ $=XX^{\top}w+Xv_{\epsilon}$ (SI.7.12)
$=M_{w}\begin{bmatrix}\theta_{\epsilon}a+a^{2}/\sqrt{d}+\theta_{q}^{2}\sqrt{d}\\ (\theta_{\epsilon}+a/\sqrt{d})p+Uq/\sqrt{d}\end{bmatrix}.$ (SI.7.13)
@@
$\frac{1}{\sqrt{d}}Xy\overset{(d)}{=}M_{w}h,$ (SI.7.14)
where $h$ is the random vector defined in (SI.7.4).
Lastly, we consider the term $y^{\top}y$ in $H_{Z}$. Since $y=X^{\top}w+v_{\epsilon}$,
$y^{\top}y$ $=\lVert X^{\top}w+v_{\epsilon}\rVert^{2}$ (SI.7.15)
$=\lVert X^{\top}w+\theta_{\epsilon}\sqrt{d}M_{v_{\epsilon}}e_{1}\rVert^{2}$ (SI.7.16)
$=(a+\theta_{\epsilon}\sqrt{d})^{2}+\theta_{q}^{2}d,$ (SI.7.17)
where the second equality follows from (SI.7.8) and to reach the last equality we have used the representation in (SI.7.10). To show (SI.7.1), we recall the definition of $H_{Z}$ in (17). Substituting (SI.7.6), (SI.7.14) and (SI.7.17) into (17), we are done. ∎
SI.8 The Stieltjes Transforms of Wishart Ensembles
@@
$\frac{1}{d}\operatorname{tr}F_{R}(\nu)\simeq\mathcal{M}_{\kappa}(\nu)$ (SI.8.2)
@@
$\mathcal{M}_{\kappa}(\nu)\coloneqq\frac{2}{\nu+1-1/\kappa+\left[(\nu+1-1/\kappa)^{2}+4\nu/\kappa\right]^{1/2}}.$ (SI.8.3)
@@
$\frac{1}{\mathcal{M}_{\kappa}(\nu)}=\frac{1}{1+\mathcal{M}_{\kappa}(\nu)/\kappa}+\nu.$ (SI.8.4)
@@
$\frac{1}{d}\operatorname{tr}F^{2}(\nu)\simeq-\mathcal{M}^{\prime}_{\kappa}(\nu)=\frac{\mathcal{M}^{2}_{\kappa}(\nu)}{1-\frac{\kappa\mathcal{M}_{\kappa}^{2}(\nu)}{[\kappa+\mathcal{M}_{\kappa}(\nu)]^{2}}}.$ (SI.8.5)
For the remainder of this appendix, we will further explore the self-consistent equation given by (SI.4.7). We will show that the solution $\chi_{\pi}$ and its derivative $\frac{d}{d\pi}\chi_{\pi}$, at $\pi=0$, can be characterized by the function $\mathcal{M}_{\kappa}(\nu)$ in (SI.8.3). To start, note that at $\pi=0$, the equation in (SI.4.7) can be written as
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I)^{-1}.$ (SI.8.6)
@@
$(E_{\mathrm{tr}}+\lambda(1+\chi_{0})I_{d+1})^{-1}=\begin{bmatrix}F_{R}(\nu_{0})+a^{\ast}(1+\rho)^{2}F_{R}(\nu_{0})b_{\mathrm{tr}}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&-a^{\ast}(1+\rho)F_{R}(\nu_{0})b_{\mathrm{tr}}\\ -a^{\ast}(1+\rho)b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})&a^{\ast}\end{bmatrix},$ (SI.8.7)
where $F_{R}(\cdot)$ is the function defined in (SI.8.1),
$\nu_{0}=\frac{1+\rho}{\alpha}+\lambda(1+\chi_{0})$ (SI.8.8)
@@
$\frac{1}{a^{\ast}}=(1+\rho)^{2}+\lambda(1+\chi_{0})-(1+\rho)^{2}b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}.$ (SI.8.9)
From (SI.8.7), the equation (SI.8.6) becomes
$\frac{\tau\chi_{0}}{1+\chi_{0}}=(1+1/d)-\frac{\lambda(1+\chi_{0})}{d}\operatorname{tr}F_{R}(\nu_{0})-(1+\rho)^{2}\frac{a^{\ast}\lambda(1+\chi_{0})}{d}\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}.$ (SI.8.10)
@@
$b_{\mathrm{tr}}^{\top}F_{R}(\nu_{0})b_{\mathrm{tr}}\leq 1\qquad\text{and}\qquad\lVert F_{R}(\nu_{0})b_{\mathrm{tr}}\rVert^{2}\leq\frac{1}{\nu_{0}}\leq\frac{\alpha}{1+\rho}.$ (SI.8.11)
Substituting the first inequality above into (SI.8.9) gives us
$a^{\ast}\lambda(1+\chi_{0})\leq 1.$ (SI.8.12)
Combining this estimate with the second inequality in (SI.8.11), we can conclude that the last term on the right-hand side of (SI.8.10) is negligible as $d\to\infty$. Moreover, using the asymptotic characterization given in (SI.8.2), the equation (SI.8.10) leads to
$\frac{\tau\chi_{0}}{1+\chi_{0}}\simeq 1-\lambda(1+\chi_{0})\mathcal{M}_{\kappa}(\nu_{0}).$ (SI.8.13)
@@
$\xi_{0}=\lambda(1+\chi_{0}),$ (SI.8.14)
and also recalling the definition of $\nu_{0}$ in (SI.8.8), we can further transform (SI.8.13) to
$\xi_{0}\mathcal{M}_{\kappa}\left(\frac{1+\rho}{\alpha}+\xi_{0}\right)-\frac{\tau\lambda}{\xi_{0}}\simeq 1-\tau.$ (SI.8.15)